OpenStudy (dj3strella):

1. Which of the following methods would be the easiest to use to solve 2x2 + 4x – 3 = 0? factoring isolating the x2 term and finding the square root of both sides using the quadratic formula all three methods would be easy and effective 2. Which of the following methods would be the easiest to use to solve x2 + 5x – 6 = 0? factoring isolating the x2 term and finding the square root of both sides using the quadratic formula all three methods would be easy and effective

OpenStudy (dj3strella):

@hero can you help me?

OpenStudy (phi):

I think isolating the x2 term and finding the square root of both sides works if you have something like x^2 - 16 =0 (no "x" term) but \[ 2x^2 + 4x – 3 = 0\] is more complicated than that (it has a 4x term)

hero (hero):

1. If it's a trinomial in the form ax^2 + bx + c and factorable, then you would use factoring. 2. If the trinomial is not factorable, then you would use the the quadratic formula. 3. If it has the form x^2 - b^2 and both a and b are perfect squares, then you can isolate x^2 If you have a trinomial and b^2 - 4ac is a positive perfect square, then the trinomial is factorable.

OpenStudy (dj3strella):

@ hero is 1 is factoring

OpenStudy (dj3strella):

and 2 is??

hero (hero):

I don't understand what you're asking